OpenStudy (anonymous):

240=200+30t-5t^2 find t

OpenStudy (anonymous):

240=200+30t-5t^2 240-200=30t - 5t^2 40=30t-5t^2 5t^2 - 30t + 40 = 0 divide by 5 t^2 - 6t + 8= 0 t^2 - 4t - 2t + 8 = 0 t(t-4)-2(t-4)=0 (t-2)(t-4)=0 therefore, t is either 4 or 2..

OpenStudy (anonymous):

thanks a lot!!

OpenStudy (anonymous):

:)